OpenStudy (anonymous):

a) On Wednesday, Harry sold 3 times as many candy bars as on Tuesday. On Thursday, he sold 5 times as many bars as on Wednesday. He sold a total of 152 candy bars. How many did he sell each day? a) On Wednesday, Harry sold 3 times as many candy bars as on Tuesday. On Thursday, he sold 5 times as many bars as on Wednesday. He sold a total of 152 candy bars. How many did he sell each day? @Mathematics

OpenStudy (turingtest):

#soldTue=x #soldWed=y #soldThur=z y=3x z=5y x+y+z=152 Three equation with three unknowns. Can you solve it from here?

OpenStudy (diyadiya):

let x be the candy bars sold on tuesday x+3x+5(3x)=152 x+3x+15x=152 19x=152 x=8(tuesday) 3x=24(wednesday) 15x=15(8)=120(thursday)